OpenStudy (ujjwal):

If the system of equation x-ky-z=0, kx-y-z=0, x+y-z=0 has a non-zero solution, the value of k are: a)-1,2 b)1,2 c)0,1 d)-1,1

OpenStudy (anonymous):

Start by Subtracting first and third equation...

OpenStudy (anonymous):

By SUBTRACTING FIRST and THIRD equation you will get k.. And SUBTRACTING FIRST and SECOND you will get other value of k..

mathslover (mathslover):

what r u getting

mathslover (mathslover):

yes and then equate them..

OpenStudy (ujjwal):

when K=-1, the solution is (0,0,0) and when K=1, we don't have a solution. So, none of the options seem correct to me. we have -1 in (a) and 1 in rest three options..

mathslover (mathslover):

I think we have atleast k = -1

mathslover (mathslover):

right ?

OpenStudy (anonymous):

are you not getting the D option??

OpenStudy (ujjwal):

K=-1 gives a solution but the solution is (0,0,0).. The qsn is asking for non-zero solution. And @waterineyes , when k=1 we don't have a solution at all. so, how can d be correct?

mathslover (mathslover):

k wait let me think about this

OpenStudy (anonymous):

At K = -1 I think the two lines have Infinitely many solutions because they are overlapping parallel lines..

mathslover (mathslover):

k = 1 that is not possible

OpenStudy (ujjwal):

My mistake... K=-1 doesn't give a solution either.

OpenStudy (anonymous):

Yes.. You think right..

mathslover (mathslover):

hmn !! so i will have to think more :)

OpenStudy (ujjwal):

I guess all options are incorrect!

mathslover (mathslover):

I found a site it is saying that d is right !!!

OpenStudy (anonymous):

No last option is right I think..

OpenStudy (anonymous):

Yes D is right @mathslover What is the problem in that..

OpenStudy (ujjwal):

umm.. how?

mathslover (mathslover):

Yes i proved that also !!! in the above images

OpenStudy (anonymous):

What is the problem you are facing??

OpenStudy (ujjwal):

so, when k=1 and k=-1 what are the solutions?

OpenStudy (anonymous):

For k =1, It is easy way to find @mathslover \[x - kx -ky + y = 0\] \[(x-y)(k-1) = 0\] \[k = 1\]

mathslover (mathslover):

yes i found that above !!!! just made it longer

mathslover (mathslover):

did u get this now @ujjwal

OpenStudy (ujjwal):

yeah, but what are the solutions to the equations when k=1 and k=-1. Th eqns are supposed to have a solution, right?

mathslover (mathslover):

this is your final solution that i got from internet

mathslover (mathslover):

I hope that this will lead u towards right answer and also right solution !!! @waterineyes what do you think about that image that i got from google books !! do u think that it is right at least i do !

OpenStudy (anonymous):

Yes it is.. On putting k = 1, first two lines become parallel but overlapping so it can have infinite number of solutions that is my logic..

mathslover (mathslover):

@ujjwal what is ur respond towards this answer now? any confusion now ?

OpenStudy (ujjwal):

Google book helps! so, we have infinite solutions in both cases, i.e. when k=1 and k=-1.. And that would be regarded as non zero solution.

mathslover (mathslover):

yes !!! thanks to google !!!

OpenStudy (ujjwal):

Well, what does non- zero solution exactly mean? Does it mean there are more than zero solutions? or, does it mean (0,0,0) is not a solution?

OpenStudy (ujjwal):

what do you exactly mean by non zero solution? That link doesn't help much..

mathslover (mathslover):

for example : x+1=0 , x=-1 this is non zero solution

mathslover (mathslover):

and if x+1=1 , x = 0 then there is no --> non-zero solution got it now ?

mathslover (mathslover):

got it now @ujjwal ?

mathslover (mathslover):

Note: if we have 2 non zero solutions a and b then we will write them as (a,b)

OpenStudy (ujjwal):

x+1=0 and x=-1 have infinite solutions.. so does non-zero solution mean infinite solution?

OpenStudy (ujjwal):

Or, does it simply mean solution other than (0,0)?

mathslover (mathslover):

see let me take another : x+y=0 if x = -1 , then y =1 x = 1 , then y = -1 and so on .. this has infinite solutions x+1=0 has finite solution that is -1

mathslover (mathslover):

Got it now ? @ujjwal

mathslover (mathslover):

btw thanks for that medal !!! so does it mean that u got it ?

OpenStudy (ujjwal):

I know what infinite solution is.. But just say me one thing: which is correct? a) Non-zero solution means that the solution to the equations (in x and y), whether it be finite or infinite, it should not be (0,0) b)Non-zero solution means there are infinite solutions to the equations (in terms of x and y) c)neither a nor b

mathslover (mathslover):

i think a is correct

OpenStudy (ujjwal):

But in above set of equations, in both cases when k=1 and k=-1, (0,0,0) satisfies all the equations...

OpenStudy (ujjwal):

I somehow get it.. finally!! thanks @mathslover and @waterineyes ..

mathslover (mathslover):

no problem and gr8 to know that